An unknown element D has two stable isotopes, 185D and 187D with masses of 184.9523 amu and 186.9564 amu. The average atomic mass of element D is 186.2071 amu. 

Determine the percentage abundances of the two isotopes.




1
Expert's answer
2022-05-06T07:11:02-0400

Ler x is a share of 185D and y - share of 187D. Then

x+y=1

184.9523x+186.9564y=186.2071

184.9523x+186.9564(1-x)=186.2071

x(184.9523-186.8564)=186.2071-186.9564

1.9041x=0.7493

x=0.3935=39.35%

y=1-0.3935=0.6065=60.65%